solve for x
75= 68+80+x/3

Answer by checkley77(12844)

75=68+80+x/3
X/3=75-68-80
X/3=-73
X=-73*3
X=-219 ANSWER.
PROOF:
75=68+80-219/3
75=68+80-73
75=75